OpenStudy (anonymous):

-14x^2+2x+9=0 For the following equation state the value of the desriminant and then desrcibe the nature of the solution.

OpenStudy (anonymous):

D=b^2-4ac =4+4*14*9...positive roots r real

OpenStudy (anonymous):

so it has one real solution?

OpenStudy (anonymous):

yes...roots are real if disc is positive

OpenStudy (anonymous):

thank you